OpenStudy (anonymous):

dy/dx = x/y and y(3)=3, then find y.

zepdrix (zepdrix):

This is in your calculus class middle? :o This is a separable differential equation, I guess you're getting an intro to these now huh :D

OpenStudy (anonymous):

Well, first that's a differential equation, type? separable variables, (i translate this in primitive, sorry for the bad english xD) Now, multiply both sides by y y dy/dx = x Then, integrate (both sides) on x \[\int\limits y\dfrac {dy} {dx} dx=\int\limits xdx\] Now, you'll obtain \[\dfrac {y^2} {2}+c =\dfrac {x^2} {2} +c\] Get together the constants, and \[y= \pm \sqrt {2c+x^2}\] Now, to find y you solve your IVP for y(3)=3, and remember that y is actually y(x) So, you'll have that \[3=\sqrt{2c+9}\] And c must be 0 \[y=sqrt {x^2}\]

OpenStudy (anonymous):

@zepdrix haha yessss :(

OpenStudy (anonymous):

did he do it correctly?

zepdrix (zepdrix):

Woops, combining the C's will give us a NEW arbitrary constant, not 2C. :D And I think it's probably better to not square root this one. Just leave it implicitly written and plug in the initial condition to solve for C.

OpenStudy (anonymous):

but thank you @Umangiasd :)

OpenStudy (anonymous):

Yeah, my mistake there, anyway, c=0 xD

zepdrix (zepdrix):

\[\large \frac{1}{2}y^2=\frac{1}{2}x^2+c\]From here, multiply both sides by 2 giving us,\[\large y^2=x^2+c\]Understand why the constant didn't change?

OpenStudy (anonymous):

ok. uhm.. could you give me a few minutes? i am going to have to try this myself. :p

OpenStudy (anonymous):

so, when i did it, i got c=-5 and so, would the final answer be \[y=\sqrt{x^2-5}\] ?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

How did you get -5? :D

OpenStudy (anonymous):

i did